OpenStudy (anonymous):

The barrel of a rifle has a length of 0.845 m. A bullet leaves the muzzle of a rifle with a speed of 568 m/s. What is the acceleration of the bullet while in the barrel? A bullet in a rifle barrel does not have constant acceleration, but constant acceleration is to be assumed for this problem. Answer in units of m/s 2

OpenStudy (anonymous):

What is acceleration?

OpenStudy (anonymous):

Ok, it is the rate of change of velocity

OpenStudy (anonymous):

what is the change of velocity for the bullet

OpenStudy (anonymous):

Well it starts at 568 m/s what I'm trying to find is the acceleration, or change of the bullets velocity

OpenStudy (anonymous):

It starts at rest that is zero. when it leaves the gun it is 568m/s

OpenStudy (anonymous):

SO we have initial velocity U, final velocity V we have distance travelled during acceleration S .

OpenStudy (anonymous):

Using the equations of motion \[V ^{2}=U ^{2} + 2AS\]

OpenStudy (anonymous):

We Make A subject of the formula

OpenStudy (anonymous):

alright so would it be 568^2=0^2*2(?)

OpenStudy (anonymous):

A=\[\frac{ V ^{2}-U ^{2} }{ 2S }\]

OpenStudy (anonymous):

Sub your values for V U and S to get your acceleration

OpenStudy (anonymous):

okay so V is 56862 U is 0 but what is S?

OpenStudy (anonymous):

S is the distance travelled in acceleration 0.845m